OpenStudy (anonymous):

Solve for y, then differentiate to get y' in terms of 'x' ; xy+2x+3x²= 4

OpenStudy (anonymous):

\[xy+2x+3x^2=4\] \[x(y+2+3x)=4\] \[y+2+3x=\frac{ 4 }{ x }\] \[y=\frac{ 4 }{ x }-3x-2\]

OpenStudy (anonymous):

That was solving for y. 1st step was to factor an x on the left hand side. Then the second step was divide both sides by x Then finally minus 2 & 3x from both sides.

OpenStudy (anonymous):

any questions? Before I move on to y'

OpenStudy (anonymous):

Hint: to take the derivative of the right hand side (with respect to y) you may use the power rule (I think that's what it's called).

OpenStudy (anonymous):

Ohhh okay no no questions.

OpenStudy (anonymous):

Thanks btw !

OpenStudy (anonymous):

notice that \[\frac{ 4 }{ x }=4*x ^{-1}\]

OpenStudy (anonymous):

y'=-4x^(-1-1)-3=-4x^-2-3=-4/x^2 -3

OpenStudy (anonymous):

got lazy sorry this is how it is really written \[y'=-4x ^{-1-1}-3=-4x ^{-2}-3=\frac{ -4 }{ x ^{2} }-3\]

OpenStudy (anonymous):

Ohh okay !! thanks

OpenStudy (anonymous):

So basically im finding y'

OpenStudy (anonymous):

@mashe im stuck please help

OpenStudy (anonymous):

Sorry burhan101 I logged off shortly after posting...I didn't neglect you on purpose...do you still need help?